Based on the table below
Quantity Price Total Cost
1 60 18
2 58 20
3 54 24
4 48 30
5 40 38
6 28 50
7 14 70

i. Calculate MC and MR
ii. Determine the quantity produced by the firm
iii. How much does the firm earn in terms
iv. What is the average cost of the firm?
v. Sketch the AC, AR, MR and MC curve without a scale.
1
Expert's answer
2019-11-01T10:27:41-0400

Q P TC TR MR MC

1 60 18 60 60 18

2 58 20 118 58 2

3 54 24 162 44 4

4 48 30 184 22 6

5 40 38 200 16 8

6 28 50 168 -32 12

7 14 70 98 -70 20


i. Calculate MC and MR

ii. The quantity produced by the firm is optimal, when MR = MC. They are equal between 5 and 6 units produced. So, the optimal quantity is Q = 5 units.

iii. The firm's profits at Q = 5 is TP = TR - TC = 200 - 38 = 162.

iv. The average cost of the firm is ATC = TC/Q.

If Q = 5, then ATC = 38/5 = 7.6.

v. AC, AR, MR are downward-sloping curves, and MC curve is upward sloping after Q = 1.